Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-01-28

Observing: Nancy Elias-Rosa, Giacomo Terreran, Aleksandar Cikota

Support Team: Stefano Benetti, Lina Tomasella

TAT: Joe Anderson, Morgan Fraser


------------------------------------------------


Weather conditions


Start of night: thin clouds passing, ~62% humidity, seeing ~1.2"

Middle of night: clear skies, good seeing (~0.7"), ~45% humidity

End of night: clear skies


------------------------------------------------


Calibrations


Evening: PESSTO_bias, PESSTO_calib, PESSTO_Sky_FF_eve_UBVRi

Morning: -


------------------------------------------------


Science


-00:30 UT: Image analysis
-00:47 UT: Follow-up: ASASSN-15uo
-01:55 UT: Follow-up: OGLE15qz (failed)
        - problem with the rotator
-02:10 UT: the EFOSC2 rotator is blocked
, known problem, restart takes ~40min
        - considering to switch to SOFI, but there are no targets in this part of night
        - 02:30 UT: problem fixed
-02:35 UT: Follow-up: OGLE15xl, seeing ~0.7"
-05:00 UT: Classification: PS16yr, seeing ~0.9"; GELATO: Type II ~ 3 months after explosion, z~0.036

-05:50 UT: Classification: OGLE16ahz; GELATO: Type Ia around maximum, z~0.1
-06:43 UT: Standard: L745a
-07:05 UT: Classification: PS16xu; GELATO: Type Ia around maximum, z~0.09
-07:41 UT: Classification: PS16uh; GELATO: Type II, >250 after explosion, z~0.025
-08:37 UT: Follow-up: PSNJ12080010, seeing ~0.9"
-09:07 UT: Standard: EG274
-09:19 UT: End of night
------------------------------------------------
Comments